OpenStudy (anonymous):

Find the missing diagonal of the rhombus that has an area of 12 feet sqaures and one diagonal of 6ft.

OpenStudy (anonymous):

@ghazi Help me ik your good at math

OpenStudy (ghazi):

do you know area of rhombus?

OpenStudy (anonymous):

we got thee answer to be 18 is that right?

OpenStudy (anonymous):

@Loujoelou

OpenStudy (anonymous):

me and him^

OpenStudy (anonymous):

we got the answer to be 18 and we couldnt find how it was wrong

OpenStudy (anonymous):

area of a rhombus = 1/2 so ... 12 = 1/2 (6 * d1) 12 = 3d1 d1 = 4 ft

OpenStudy (ghazi):

it's 1/2*x*y=12.........1/2*6*y=12......y= 12/3=4 is the answer

OpenStudy (anonymous):

ok thanks

OpenStudy (ghazi):

:)